OpenStudy (anonymous):

Find the points at which the tangent line is horizontal.. f(x) = cos x + sin x; 0 <=x <= 2x

jimthompson5910 (jim_thompson5910):

f(x) = cos x + sin x f'(x) = -sin x + cos x Now set the derivative function to zero and solve for x f'(x) = -sin x + cos x 0 = -sin x + cos x sin x = cos x Use the unit circle to see that the solution is x = pi/4 So the tangent line is horizontal at x = pi/4 Note: You can add multiples of pi to the solution for x to find other solutions.

OpenStudy (anonymous):

Hi jim_thompson5190, Thanks for your answer. That's what I got too, but the back of the book gives the following answer. (pi/4, sqrt(2)), (5pi/4, - sqrt(2)) That has me a bit confused...

jimthompson5910 (jim_thompson5910):

pi/4 is the x value and f(pi/4) = sqrt(2) (ie plug in x = pi/4 into the original function and simplify). So one point where the tangent is horizontal is (pi/4, sqrt(2)) Now add pi to pi/4 to get 5pi/4. This is another value of x where the tangent is horizontal. It turns out that f(5pi/4) = -sqrt(2), which means that another point where the the tangent is horizontal is (5pi/4, -sqrt(2))

OpenStudy (anonymous):

Thanks so much for your answer!